OpenStudy (anonymous):

Let A and B be 2 x 2 matrices with integer entries such that A, A + B, A + 2B, A + 3B, and A + 4B are all invertible matrices whose inverses have integer entries. Show that A + 5B is invertible and that its inverse has integer entries.

OpenStudy (anonymous):

This is a problem in which i have no idea where to start, just looking for some ideas or suggestions :)

OpenStudy (anonymous):

What? Something you don't know. I'm surprised to see

OpenStudy (anonymous):

If you don't know, I don't think anyone else knows

OpenStudy (anonymous):

i would like alchemista to look at it, he/she is pro:) im not really looking for a solution, just ideas.

OpenStudy (anonymous):

I don't think alchemista is around at the moment

OpenStudy (anonymous):

i can tell :( im gonna mess with it for a bit.

OpenStudy (anonymous):

joemath can u tell me abt invertible matrices... i forgot abt them

OpenStudy (anonymous):

well, there are a couple of things, namely that the det of the matrix isnt 0, or that there exists another matrix such that when you multiply them you get the identity matrix.

OpenStudy (anonymous):

what if adding the matrices helps us...

OpenStudy (anonymous):

we can have a new matrice if we add a+3b and a+2b

OpenStudy (anonymous):

and subtracting a

OpenStudy (anonymous):

woul still give us an invertible matrice and with int in it

OpenStudy (anonymous):

so you have that A+5B = (A+2B) + (A + 3B) - (A)

OpenStudy (anonymous):

how do you know that? i dont think that is necessarily true.

OpenStudy (anonymous):

a have some int so if we subtract a frm 2a it still would be int

OpenStudy (anonymous):

thats right, we know that A + 5B is going to have integers in it, but we dont know that its inverse is

OpenStudy (anonymous):

The inverse of a matrix will have integer entries if the determinant of the matrix is 1 or -1

OpenStudy (anonymous):

is that an "if and only if" statement?

OpenStudy (anonymous):

yes of course because

OpenStudy (anonymous):

if the det is nonzero its invertible

OpenStudy (anonymous):

err wait a second

OpenStudy (anonymous):

determinants of matrice is non zero then if we add two non zero det matrice don't u think the result would be non zero only ? i'm not sure abt tht but i've an intuition fr it

OpenStudy (anonymous):

the reverse implication is, if the inverse has integer entries is the determinant of the original -1 or 1

OpenStudy (anonymous):

The intuition is based on elementary operations again

OpenStudy (anonymous):

I dont know if it will help you with the problem.

OpenStudy (anonymous):

anything is helpful :) although im gonna go to bed, thinking about it now is making me tired lol.

OpenStudy (anonymous):

i'll look at it from that angle tomorrow when i have time, thank you both for your input :)

OpenStudy (anonymous):

Also make sure you are thinking about this \[\det \left[\begin{matrix}a & b \\ c & d\end{matrix}\right] = ad - bc\]

OpenStudy (anonymous):

yeah, i have that written out for both A and B in an arbitrary way, but there are so many variables its a tad confusing.

OpenStudy (anonymous):

anywhos, sleeeeeep, thanks again.

OpenStudy (anonymous):

a and b must be nonzero ....

OpenStudy (anonymous):

so 4b will only give 4det b

OpenStudy (anonymous):

which will be nonzero

OpenStudy (anonymous):

or even if b is zero then a can't be zero.... so a + 5b must be non zero matrice

OpenStudy (anonymous):

alchemista in the question its already given that a+4b is non zero...so either a is non zero or b or maybe both..........then a + 5 b should be non zero too...

OpenStudy (anonymous):

we aren't talking about weather the matrix has non zero entries.

OpenStudy (anonymous):

invertibility is a stronger condition

OpenStudy (anonymous):

there are nonzero 2x2 matrices that are not invertible

OpenStudy (anonymous):

when i say non zero i mean non zero determinant

OpenStudy (anonymous):

yeah but its given in the question that a is invertible

OpenStudy (anonymous):

if we can sum a+3b then we can divide them too (A) + 3(B)

OpenStudy (anonymous):

and it doesn't matter whether b is invertible or not .....becuz we know a is..

OpenStudy (anonymous):

You will need to show the following \[(a_1+ka_2)(d_1+kd_2)-(b_1+kb_2)(c_1+kc_2)\] is 1 or -1

OpenStudy (anonymous):

for k = 5 given that it holds for 1,2 ,3, 4

OpenStudy (anonymous):

You are also given that \[a_1d_1 - b_1c_1\] is 1 or -1

OpenStudy (anonymous):

now we know a1d1 - b1c1 /= 0

OpenStudy (anonymous):

from det of a

OpenStudy (anonymous):

What I said before is an if and only if. Since the inverse has integer entries the det is 1/-1

OpenStudy (anonymous):

1 or -1